OpenStudy (anonymous):

The heat developed in an electric wire varies jointly as the wire's resistance, the time the current flows, and the square of the current. In two minutes a current of 5 amps develops 1,200 heat units in a wire of 8 ohms resistance. What resistance does a similar wire have, which develops 6,000 heat units with a current of 10 amps in 5 minutes?

OpenStudy (wasiqss):

4 ohm

OpenStudy (anonymous):

h = k*r*t*i^2 -> heat = k*resistance*time*current^2 1200 = k*8*2*25 6000 = k*r*5*100 k*8*2*25/1200 = k*r*5*100/6000 r = (6000*16*25)/(1200*500) = 0.8